OpenStudy (anonymous):

Find volume of solid bounded by: cylinder (x^2+y^2=a^2), surface (z=x^2) and plane (z=0) I wanted to find the volume by integrating int[region] dxdydz The question is, how do I describe the region in that configuration?

OpenStudy (anonymous):

We have : \[\Large V=\int_Ddxdydz\] Where : \[\Large D=\{(x,y,z)\in\mathbb R^3, x^2+y^2\leq a^2,~ 0\leq z\leq x^2~~\}\] As I think !